OpenStudy (yumeybitaa):

Jack earns $50 for washing cars on the weekend. He makes an average of $4.25 in tips per hour. Write the function of Joe's earnings, and solve for how much he would make after 6 hours. y = 4.25x + 50; $54.25 y = 50x + 4.25; $69.50 y = 50x + 4.25; $304.25 y = 4.25x + 50; $75.50

OpenStudy (anonymous):

It is actually the last one because: Y = 4.25(6) + 50 4.25 x 6 = 25.5 25.5 + 50 = 75.50

OpenStudy (johnweldon1993):

He makes a constant of 50 dollars, plus a variable of 4.25h where 'h' how may hours he works \[\large y = 4.25h + 50\] Now for working 6 hours we have \[\large y = 4.25(6) + 50\] \[\large y = 25.50 + 50\] \[\large y = 75.50\]

OpenStudy (anonymous):

Yes that is what I was going to say! The last one is correct because you have multiply the tips and then add it to the 50$ he made over the weekend.

OpenStudy (anonymous):

can you guys help me on my problem

OpenStudy (yumeybitaa):

Thank you :)